# FreeScale: Unleashing the Resolution of Diffusion Models via Tuning-Free Scale Fusion

### v1

@@ Appendix C Gallery of 8k Images
Appendix D Limitation and Future Work
@@
Generalization. DiT-based LDMs (e.g., FLUX [29] and CogVideoX [48]), have showcased impressive visual generation capabilities recently. When DiT-based LDMs are required to directly generate higher-resolution results, obvious quality degradation also occurs, but in a significantly different way than in the UNet-based LDMs. Therefore, FreeScale currently only supports the UNet-based LDMs without further modification. One temporary solution is using DiT-based LDMs to generate the intermediate 1×1$\times$1 × result and then utilizing the remaining pipeline of FreeScale to generate higher-resolution results. In this way, we successfully obtain the dragon with “FreeScale” in Figure 8, leveraging FLUX’s superior visual text rendering capabilities. However, this approach results in a slight loss of quality due to the gap in understanding of the same text by different models. The final solution still needs to be customized specifically for DiT-based LDMs.

### v2

@@ C.3 Gallery of 8k Images
Appendix D Limitations and Future Work
@@
Structure Gap. DiT-based LDMs (e.g., FLUX [33] and CogVideoX [55]), have showcased impressive visual generation capabilities recently. However, UNet-based LDMs and DiT-based LDMs will face different challenges in the higher-resolution generation task. As shown in Figure 13, UNet-based LDMs face repetition problems while DiT-based LDMs face blur problems. Most previous higher-resolution generation methods either support the UNet-based LDMs (DemoFusion [14], and FouriScale [25]) or DiT-based LDMs (I-MAX [15]), in line with the common sense that different problems require different strategies to solve. To support the DiT based structure, FreeScale also needs to be customized specifically.